OpenStudy (anonymous):

How do you calculate 0.48351=(0.94495)^n

OpenStudy (anonymous):

use logarithms

OpenStudy (anonymous):

Yep, logarithm will do. log(0.48351)=log(0.94495)^n log(0.48351)=n*log(0.94495) n= ................... (find it yourself :))

OpenStudy (anonymous):

or more simply put ... \[n=\log_{0.48351}0.94495 \]

OpenStudy (anonymous):

@aceace : do your browser convert your latex code? Strangely, mine don't....

OpenStudy (anonymous):

yeh but you have to wait a while... like 20 seconds

OpenStudy (anonymous):

It works now. I just restarted my browser. Didn't know what happened but it works :D